# Controlling the Flow: Stability and Convergence for Stochastic Gradient Descent with Decaying Regularization

### v1

@@ 1 Introduction
In this work we study the unconstrained optimization problem
@@
where $(\mathcal{X},\langle\cdot,\cdot\rangle_{\mathcal{X}})$ is a separable real Hilbert space with inner product $\langle\cdot,\cdot\rangle_{\mathcal{X}}$ and induced norm $\|x\|^{2}_{\mathcal{X}}=\langle x,x\rangle_{\mathcal{X}}$. The objective function $f:\mathcal{X}\to\mathbb{R}$ will be assumed to be differentiable, convex, $L$-smooth with $\operatorname{arg\min}_{x\in\mathcal{X}}f(x)\neq\emptyset$. Moreover, we will always denote by $x_{\ast}\in\operatorname{arg\min}_{x\in\mathcal{X}}f(x)$ the minimum-norm solution, i.e. a minimum with $\|x_{\ast}\|_{\mathcal{X}}\leq\|\hat{x}\|_{\mathcal{X}}$ for all $\hat{x}\in\operatorname{arg\min}_{x\in\mathcal{X}}f(x)$. A common strategy for finding a point close to the minimum-norm solution is to employ regularization techniques. One popular approach from the optimization literature is to include Tikhonov regularization into (1) in the form of
@@
where $\lambda\geq 0$ is called the regularization parameter. Since the regularized objective function $f_{\lambda}$ is $\lambda$-strongly convex for any $\lambda>0$, there exists a unique minimum $x_{\lambda}=\operatorname{arg\min}_{x\in\mathcal{X}}f_{\lambda}(x)$ and many first order methods, such as stochastic gradient descent (SGD), are able to efficiently find $x_{\lambda}$.
Tikhonov regularization is a simple but effective method that appears in various contexts, such as statistics (e.g. ridge regression, Hoerl ), classical inverse problems EHN1996 , including parameter estimation in partial differential equations isakov2006inverse and image reconstruction ito2014inverse ; bertero2021introduction , dating all the way back to Tikhonov tikhonov1943stability . In the context of training neural networks, Tikhonov regularization is known under the name weight decay as the method decreases the norm of the neural network weights. One early reference is Krogh , for more recent work on the effect of weight decay on generalization we refer to Shalev , for LLM training to Brown , and for a very recent experimental deep learning study to Angelo . It is still a very much open problem to fully understand the different effects on weight decay, both from a practical but also the theoretical point of view in different optimization settings.
@@ 1.1 Fixing the setup
which for constant $\lambda$ converges to $x_{\lambda}$ under suitable conditions on the step-size sequence $\alpha$. In many applications the gradient cannot be computed (or observed) exactly, instead only gradients with noisy perturbation are available. This leads to two equivalent formulations: one in which a noisy perturbation $D_{k}$ is added to the true gradient, and another in which the gradient is replaced by an estimated gradient $\widehat{\nabla f(X_{k-1})}$. These formulations are equivalent if we define the noise as $D_{k}=\widehat{\nabla f(X_{k-1})}-\nabla f(X_{k-1})$. We thus stick to the first setting but use the more accessible second notation for the pseudocode of Algorithm 1 below.
@@
where $\mathbb{E}[\|X_{0}\|_{\mathcal{X}}^{2}]<\infty$, $\alpha$ and $\lambda$ are sequences of (deterministic or random) non-negative reals, and $D:=(D_{k})_{k\in\mathbb{N}}$ is an adapted sequence of martingale differences, i.e. $\mathbb{E}[D_{k}\mid\mathcal{F}_{k-1}]=0$ for all $k\in\mathbb{N}$. More precisely, in Theorem 2.1 we assume the sequences $\alpha:=(\alpha_{k})_{k\in\mathbb{N}}$ and $\lambda:=(\lambda_{k})_{k\in\mathbb{N}}$ to be predictable stochastic processes, i.e. $\alpha_{k}$ and $\lambda_{k}$ are $\mathcal{F}_{k-1}$-measurable for all $k\in\mathbb{N}$. The SGD formalism includes for instance stochastic gradients in finite-sum problems, where a random data point’s gradient estimates the full gradient, and in expected risk minimization, where gradients are computed using samples from the data distribution.
@@
0:  Initial guess $X_{0}$, number of iterations $N$, step-size schedule $\alpha$, regularization schedule $\lambda$
1:  for $k=1$ to $N$ do
@@
The objective function $f:\mathcal{X}\to\mathbb{R}$ is convex, continuously differentiable, and $L$-smooth. The latter means that $\nabla f:\mathcal{X}\to\mathcal{X}$ is globally $L$-Lipschitz continuous, i.e. there exists $L>0$ such that $\|\nabla f(x)-\nabla f(y)\|_{\mathcal{X}}\leq L\|x-y\|_{\mathcal{X}}$ for all $x,y\in\mathcal{X}$. Furthermore, we assume that $\operatorname{arg\min}_{x\in\mathcal{X}}f(x)\neq\emptyset$ and denote by $x_{\ast}\in\operatorname{arg\min}_{x\in\mathcal{X}}f(x)$ the minimum-norm solution.
@@
$\mathbb{E}\big{[}\|D_{k}\|_{\mathcal{X}}^{2}\big{|}\mathcal{F}_{k-1}\big{]}\leq A(f(X_{k-1})-f(x_{\ast}))+C,\quad k\in\mathbb{N}.$
@@
see e.g. Lemma C.1 in Wojtowytsch2023 . The exact same argument (combining $L$-smoothness and the fundamental theorem of calculus) extends readily to the Hilbert space setting. Therefore, 1.2 is equivalent to the classical ABC-condition
$\mathbb{E}\big{[}\|D_{k}\|_{\mathcal{X}}^{2}\big{|}\mathcal{F}_{k-1}\big{]}\leq A(f(X_{k-1})-f(x_{\ast}))+B\|\nabla f(X_{k-1})\|_{\mathcal{X}}^{2}+C,\quad k\in\mathbb{N},$
@@ 1.2 Contribution
appearing on the right-hand side. The main achievement of this article is to carry out last-iterate estimates that yield $L^{2}$ and almost sure convergence rates. In contrast to non-regularized SGD we obtain convergence to the minimum-norm solution (not just some solution) of the optimization problem while obtaining comparable rates for the optimality gap. The simulation on the right shows the effect for $f(x_{1},x_{2})=(x_{1}+x_{2}-1)^{2}$ with noisy gradients perturbed by independent Gaussians. While vanilla SGD converges to some minima (red dots), reg-SGD converges to the minimum-norm solution. Another important theoretical property that we reveal is that reg-SGD is more stable. It turns out that the iterative scheme is automatically bounded and cannot explode.
@@ 1.3 Related work
Deterministic Tikhonov regularization in continuous time: The literature contains a number of articles on Tikhonov regularization with decreasing regularization parameter, mostly in the context of deterministic optimization in continuous time. This includes the analysis of first order ODEs attouch1996dynamical ; cominetti2008strong ; attouch2022acceleratedgradientmethodsstrong and second order ODEs jendoubi2010asymptotically ; attouch2018combining ; alecsa2021tikhonov ; attouch2022acceleratedgradientmethodsstrong ; attouch2024convex . Extensions to stochastic optimization in continuous time, in particular an analysis of the stochastic differential inclusion process, have been considered in maulen2024tikhonov . Many statements are based on results for the solution curve $(x_{\lambda})_{\lambda\geq 0}$ derived in attouch1996viscosity . In this article, we extend some of their ideas for steepest descent flows to the discrete-time (and stochastic) setting. To the best of our knowledge, the only prior work on decaying Tikhonov regularization in discrete time algorithms is laszlo2023convergence , which analyzes the convergence of a deterministic inertial proximal algorithm with a polynomially decaying regularization parameter.
@@
Stochastic gradient descent (with regularization): For recent results on convergence of SGD for possibly non-convex optimization landscapes we refer the reader to mertikopoulos2020sure ; dereich2021convergence ; weissmann2025almost and references therein. Note that due to the non-convexity, only convergence to a critical point can be shown without guarantees of optimality. In the smooth and convex case, almost sure convergence rates for the last-iterates of SGD without Tikhonov regularization under the ABC-condition for the noise have been derived in JMLR:v25:23-1436 . Therein, it was proved that for step-sizes $\alpha_{k}=C_{\alpha}k^{-\frac{2}{3}-\varepsilon}$ with $\varepsilon\in(0,\frac{1}{3})$ one has $f(X_{k})-f(x_{\ast})\in\mathcal{O}(k^{-\frac{1}{3}+\varepsilon})$ almost surely. moulines2011non gives a rate of convergence for SGD with polynomially decaying step-sizes in expectation. Their rate is optimal for the choice $\alpha_{k}=C_{\alpha}k^{-\frac{2}{3}}$ which yield the rate $\mathbb{E}[f(X_{k})-f(x_{\ast})]\in\mathcal{O}(k^{-1/3})$. Assuming uniform boundedness of the gradients and iterates, shamir2013stochastic increased this rate of convergence to $\mathcal{O}(k^{-1/2}\log(k))$ for the step-sizes $\alpha_{k}=C_{\alpha}k^{-\frac{1}{2}}$. These additional assumptions have been lifted in liu2024revisiting , where only bounded variance of the noise is assumed. Following moulines2011non ; liu2024revisiting , the Ruppert-Polyak average also achieves the rate $\mathbb{E}[f(\bar{X}_{k})-f(x_{\ast})]\in\mathcal{O}(k^{-1/2}\log(k))$ for $\alpha_{k}=C_{\alpha}k^{-\frac{1}{2}}$. These results attain the lower bound for the optimization of $L$-smooth convex functions using first order algorithms that have access to unbiased gradient estimates with bounded variance derived in agarwal2009information up to a $\log(k)$ factor. More recently, almost sure convergence rates under a related setting have been derived in pmlr-v134-sebbouh21a . However, without any additional regularization one can only guarantee convergence towards some global minimum. The present article targets specifically algorithms that find the minimum-norm solution. We show that, when using reg-SGD one gets comparable convergence rates in the optimality gap as the ones for vanilla SGD cited above. Moreover, one can weaken the bounded variance assumption on the noise, while also achieving strong convergence to the minimum-norm solution. Finally, we point to HE20181 , where the role of regularization for the convergence of SGD for a prescribed number of optimization steps is discussed. Using a fixed regularization parameter, the authors derive a complexity bound for averaged SGD. However, they do not discuss the role of Tikhonov regularization for convergence towards a minimum-norm solution.
@@ 2 Theoretical results
In this section, we present our main theoretical contributions concerning the convergence of stochastic gradient descent with decaying Tikhonov regularization. An abstract convergence result is presented, followed by quantitative rates of convergence for polynomial step-size and regularization schedules. For the ML practitioner, we derive optimal choices for the step-sizes and regularization parameters. All proofs are provided in Appendix C, where slightly more general statements are presented.
Approach: First, we carefully balance the step-size and the regularization parameters to ensure convergence of the energy function $E_{k}=f_{\lambda_{k+1}}(X_{k})-f_{\lambda_{k+1}}(x_{\lambda_{k+1}}).$ In Lemma B.3, we then obtain the estimate $\|X_{k}-x_{\lambda_{k}}\|_{\mathcal{X}}\leq E_{k}/\lambda_{k},$ which links the distance to the regularized minimizer with the energy function. Since $\lambda_{k}$ also influences the decay of $E_{k}$, we must jointly control both quantities to ensure that $\lim_{k\to\infty}E_{k}/\lambda_{k}=0.$ Combined with the fact that $\lim_{\lambda\to 0}\|x_{\lambda}-x_{\ast}\|_{\mathcal{X}}=0,$ this yields strong convergence $\lim_{k\to\infty}\|X_{k}-x_{\ast}\|_{\mathcal{X}}=0$ (both in $L^{2}$ and almost surely). Moreover, if a convergence rate for $\|x_{\lambda}-x_{\ast}\|_{\mathcal{X}}$ is known, the analysis allows us to also quantify a strong convergence rate for the iterates $X_{k}$ to $x_{\ast}$.

### v2

@@ 1 Introduction
In00footnotetext: These authors contributed equally to this work. this work we study the unconstrained optimization problem
@@
where $(\mathcal{X},\langle\cdot,\cdot\rangle_{\mathcal{X}})$ is a separable real Hilbert space with inner product $\langle\cdot,\cdot\rangle_{\mathcal{X}}$ and induced norm $\|x\|^{2}_{\mathcal{X}}=\langle x,x\rangle_{\mathcal{X}}$. The objective function $f:\mathcal{X}\to\mathbb{R}$ will be assumed to be differentiable, convex, and $L$-smooth with $\operatorname*{arg\,min}_{x\in\mathcal{X}}f(x)\neq\emptyset$. Moreover, we will always denote by $x_{\ast}\in\operatorname*{arg\,min}_{x\in\mathcal{X}}f(x)$ the minimum-norm solution, i.e. a minimum with $\|x_{\ast}\|_{\mathcal{X}}\leq\|\hat{x}\|_{\mathcal{X}}$ for all $\hat{x}\in\operatorname*{arg\,min}_{x\in\mathcal{X}}f(x)$. A common strategy for finding a point close to the minimum-norm solution is to employ regularization techniques. One popular approach from the optimization literature is to include Tikhonov regularization into (1) in the form of
@@
where $\lambda\geq 0$ is called the regularization parameter. Since the regularized objective function $f_{\lambda}$ is $\lambda$-strongly convex for any $\lambda>0$, there exists a unique minimum $x_{\lambda}=\operatorname*{arg\,min}_{x\in\mathcal{X}}f_{\lambda}(x)$ and many first order methods, such as stochastic gradient descent (SGD), are able to efficiently find $x_{\lambda}$.
Tikhonov regularization is a simple but effective method that appears in various contexts, such as statistics (e.g. ridge regression, Hoerl ), classical inverse problems EHN1996 , including parameter estimation in partial differential equations isakov2006inverse and image reconstruction ito2014inverse ; bertero2021introduction , dating all the way back to Tikhonov tikhonov1943stability . In the context of training neural networks, Tikhonov regularization is known under the name weight decay as the method decreases the norm of the neural network weights. One early reference is Krogh , for more recent work on the effect of weight decay on generalization we refer to Shalev , for LLM training to Brown , and for a very recent experimental deep learning study to Angelo . It is still a very much open problem to fully understand the different effects of weight decay, both from a practical but also the theoretical point of view in different optimization settings.
@@ 1.1 Fixing the setup
which for constant $\lambda$ converges to $x_{\lambda}$ under suitable conditions on the step-size sequence $\alpha$. In many applications the gradient cannot be computed (or observed) exactly, instead only gradients with noisy perturbation are available. This leads to two equivalent formulations: one in which a noisy perturbation $D_{k}$ is added to the true gradient, and another in which the gradient is replaced by an estimated gradient $\widehat{\nabla f(X_{k-1})}$. These formulations are equivalent if we define the noise as $D_{k}=\widehat{\nabla f(X_{k-1})}-\nabla f(X_{k-1})$. We thus stick to the first setting but use the more accessible second notation for the pseudocode of Algorithm˜1 below.
@@
where $\mathbb{E}[\|X_{0}\|_{\mathcal{X}}^{2}]<\infty$, $\alpha$ and $\lambda$ are sequences of (deterministic or random) non-negative reals, and $D:=(D_{k})_{k\in\mathbb{N}}$ is an adapted sequence of martingale differences, i.e. $\mathbb{E}[D_{k}\mid\mathcal{F}_{k-1}]=0$ for all $k\in\mathbb{N}$. More precisely, in Theorem˜2.1 we assume the sequences $\alpha:=(\alpha_{k})_{k\in\mathbb{N}}$ and $\lambda:=(\lambda_{k})_{k\in\mathbb{N}}$ to be predictable stochastic processes, i.e. $\alpha_{k}$ and $\lambda_{k}$ are $\mathcal{F}_{k-1}$-measurable for all $k\in\mathbb{N}$. The SGD formalism includes for instance stochastic gradients in finite-sum problems, where a random data point’s gradient estimates the full gradient, see Section˜1.1 below, and in expected risk minimization, where gradients are computed using samples from the data distribution.
@@
0: Initial guess $X_{0}$, number of iterations NN, step-size schedule $\alpha$, regularization schedule $\lambda$
1:for $k=1$ to NN do
@@
The objective function $f:\mathcal{X}\to\mathbb{R}$ is convex, continuously differentiable, and $L$-smooth. The latter means that $\nabla f:\mathcal{X}\to\mathcal{X}$ is globally $L$-Lipschitz continuous, i.e. there exists $L>0$ such that $\|\nabla f(x)-\nabla f(y)\|_{\mathcal{X}}\leq L\|x-y\|_{\mathcal{X}}$ for all $x,y\in\mathcal{X}$. Furthermore, we assume that $\operatorname*{arg\,min}_{x\in\mathcal{X}}f(x)\neq\emptyset$ and denote by $x_{\ast}\in\operatorname*{arg\,min}_{x\in\mathcal{X}}f(x)$ the minimum-norm solution.
@@
$\mathbb{E}\big[\|D_{k}\|_{\mathcal{X}}^{2}\mid\mathcal{F}_{k-1}\big]\leq A(f(X_{k-1})-f(x_{\ast}))+C,\quad k\in\mathbb{N}.$
@@
see e.g. Lemma C.1 in Wojtowytsch2023 . The exact same argument (combining $L$-smoothness and the fundamental theorem of calculus) extends readily to the Hilbert space setting. Therefore, Section˜1.1 is equivalent to the classical ABC-condition
$\mathbb{E}\big[\|D_{k}\|_{\mathcal{X}}^{2}\mid\mathcal{F}_{k-1}\big]\leq A(f(X_{k-1})-f(x_{\ast}))+B\|\nabla f(X_{k-1})\|_{\mathcal{X}}^{2}+C,\quad k\in\mathbb{N},$
@@
Example 1.3 (Mini-batch estimator for finite-sum problems).
Consider the finite-sum optimization problem
$\min_{x\in\mathbb{R}^{d}}f(x)=\frac{1}{N}\sum_{i=1}^{N}f_{i}(x),$
where, for all $i=1,\dots,N$, $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ is convex and $L_{i}$-smooth. At iteration $k\in\mathbb{N}$, we can define a mini-batch estimator with mini-batch size $M\in\mathbb{N}$ via $g_{k}=\frac{1}{M}\sum_{i\in M}\nabla f_{I_{i,k}}(X_{k-1}),$ where $(I_{i,k})_{i,k\in\mathbb{N}}$ is a family of iid. random variables that are uniformly distributed on $\{1,\dots,N\}$. The corresponding gradient noise is defined as $D_{k}=\frac{1}{M}\sum_{i\in M}(\nabla f_{I_{i,k}}(X_{k-1})-\nabla f(X_{k-1}))$ and satisfies
$\mathbb{E}\left[\|D_{k}\|^{2}\mid\mathcal{F}_{k-1}\right]\leq\frac{4L}{M}\bigl(f(X_{k-1})-f(x_{\ast})\bigr)+\frac{2\sigma_{\ast}^{2}}{M},$ (6)
where $\bar{L}=\frac{1}{N}\sum_{i=1}^{N}L_{i}$ and $\sigma_{\ast}^{2}=\frac{1}{N}\sum_{i=1}^{N}\|\nabla f_{i}(x_{\ast})\|^{2}$. We will prove (6) in Appendix˜C below.
@@ 1.2 Contribution
appearing on the right-hand side. The main achievement of this article is to carry out last-iterate estimates that yield $L^{2}$ and almost sure convergence rates. In contrast to non-regularized SGD we obtain convergence to the minimum-norm solution (not just some solution) of the optimization problem while obtaining comparable rates for the optimality gap. The simulation in Figure˜1 on the right shows the effect for $f(x_{1},x_{2})=(x_{1}+x_{2}-1)^{2}$ with noisy gradients perturbed by independent Gaussians. While vanilla SGD converges to some minima (red dots), reg-SGD converges to the minimum-norm solution. Another important theoretical property that we reveal is that reg-SGD is more stable. It turns out that the iterative scheme is automatically bounded and cannot explode.
@@ 1.3 Related work
Deterministic Tikhonov regularization: The literature contains a number of articles on Tikhonov regularization with decreasing regularization parameter. For instance, in the context of deterministic optimization, this includes the analysis of first order ODEs attouch1996dynamical ; cominetti2008strong ; attouch2022acceleratedgradientmethodsstrong and second order ODEs jendoubi2010asymptotically ; attouch2018combining ; alecsa2021tikhonov ; attouch2022acceleratedgradientmethodsstrong ; attouch2024convex . Extensions to stochastic optimization in continuous time, in particular an analysis of the stochastic differential inclusion process, have been considered in maulen2024tikhonov . Many statements are based on results for the solution curve $(x_{\lambda})_{\lambda\geq 0}$ derived in attouch1996viscosity . More generally, differential inclusions for constrained convex optimization problems have been intensively studied in ATTOUCH20101315 ; ACP2011 ; attouch2011prox ; czarnecki2016splitting ; peypouquet_coupling_2012 . In Section˜D.2, we provide more details and illustrate the relation to Tikhonov regularization. Recently, the gradient flow for a fixed Tikhonov regularization has been analyzed in boursier2025theoretical . For small values of the regularization parameter $\lambda$, the optimization dynamics can be decomposed into two distinct phases: an initial fast convergence toward the set of minima, followed by a slow drift along this set that selects the minimizer with the smallest $\ell^{2}$-norm.
In this article, we extend methodologies for steepest descent flows to the stochastic discrete-time setting. Discrete time algorithms with decaying Tikhonov regularization have been analyzed in the context of iterative regularization schemes. For instance, in Chapter 5 of bakushinsky1994illposed iterative regularization is discussed to solve variational inequalities covering (3) for convex $f$ as a special case. Under certain conditions on the regularization decay and the step-sizes, strong convergence to the minimum-norm solution can be guaranteed. A related analysis has been considered for non-linear inverse problems bakushinsky2004iterative . In the specific application to inverse problems, (3) is also known as the modified Landweber iteration, where convergence is mainly studied for nonlinear forward models using a-priori and a-posteriori stopping rules bakushinsky2004iterative ; KaltenbacherNeubauerScherzer2008 ; scherzer1998modified . The theoretical analysis is conducted in a non-convex setting and relies on the so-called tangential cone condition.
@@
Stochastic gradient descent (with regularization): For recent results on convergence of SGD for possibly non-convex optimization landscapes we refer the reader to mertikopoulos2020sure ; dereich2021convergence ; weissmann2025almost and references therein. Note that due to the non-convexity, only convergence to a critical point can be shown without guarantees of optimality. In the smooth and convex case, almost sure convergence rates for the last-iterates of SGD without Tikhonov regularization under the ABC-condition for the noise have been derived in JMLR:v25:23-1436 . Therein, it was proved that for step-sizes $\alpha_{k}=C_{\alpha}k^{-\frac{2}{3}-\varepsilon}$ with $\varepsilon\in(0,\frac{1}{3})$ one has $f(X_{k})-f(x_{\ast})\in\mathcal{O}(k^{-\frac{1}{3}+\varepsilon})$ almost surely. moulines2011non gives a rate of convergence for SGD with polynomially decaying step-sizes in expectation. Their rate is optimal for the choice $\alpha_{k}=C_{\alpha}k^{-\frac{2}{3}}$ which yield the rate $\mathbb{E}[f(X_{k})-f(x_{\ast})]\in\mathcal{O}(k^{-1/3})$. Assuming uniform boundedness of the gradients and iterates, shamir2013stochastic increased this rate of convergence to $\mathcal{O}(k^{-1/2}\log(k))$ for the step-sizes $\alpha_{k}=C_{\alpha}k^{-\frac{1}{2}}$. These additional assumptions have been lifted in liu2024revisiting , where only bounded variance of the noise is assumed. Following moulines2011non ; liu2024revisiting , the Ruppert-Polyak average also achieves the rate $\mathbb{E}[f(\bar{X}_{k})-f(x_{\ast})]\in\mathcal{O}(k^{-1/2}\log(k))$ for $\alpha_{k}=C_{\alpha}k^{-\frac{1}{2}}$. These results attain the lower bound for the optimization of $L$-smooth convex functions using first order algorithms that have access to unbiased gradient estimates with bounded variance derived in agarwal2009information up to a $\log(k)$ factor. More recently, almost sure convergence rates under a related setting have been derived in pmlr-v134-sebbouh21a . However, without any additional regularization one can only guarantee convergence towards some global minimum. The present article targets specifically algorithms that find the minimum-norm solution. We show that, when using reg-SGD one gets comparable convergence rates in the optimality gap as the ones for vanilla SGD cited above. Moreover, one can weaken the bounded variance assumption on the noise, while also achieving strong convergence to the minimum-norm solution. We also point to HE20181 , where the role of regularization for the convergence of SGD for a prescribed number of optimization steps is discussed. Using a fixed regularization parameter, the authors derive a complexity bound for averaged SGD. However, they do not discuss the role of Tikhonov regularization for convergence towards a minimum-norm solution. Finally, in the context of inverse problems the regularization properties of (vanilla) SGD have been analyzed for linear Jahn_2020 ; Jin_2019 ; doi:10.1137/20M1374456 and non-linear forward models doi:10.1137/19M1271798 based on a-priori and a-posteriori stopping rules. Moreover, in NEURIPS2022_3e8b1835 SGD has been considered for inverse problems from a statistical point of view.
@@ 2 Theoretical results
In this section, we present our main theoretical contributions concerning the convergence of stochastic gradient descent with decaying Tikhonov regularization. An abstract convergence result is presented, followed by quantitative rates of convergence for polynomial step-size and regularization schedules. For the ML practitioner, we derive optimal choices for the step-sizes and regularization parameters. All proofs are provided in Appendix˜D, where slightly more general statements are presented.
Approach: First, we carefully balance the step-size and the regularization parameters to ensure convergence of the energy function $E_{k}=f_{\lambda_{k+1}}(X_{k})-f_{\lambda_{k+1}}(x_{\lambda_{k+1}}).$ In Appendix˜B, we then obtain the estimate $\|X_{k}-x_{\lambda_{k}}\|_{\mathcal{X}}\leq E_{k}/\lambda_{k},$ which links the distance to the regularized minimizer with the energy function. Since $\lambda_{k}$ also influences the decay of $E_{k}$, we must jointly control both quantities to ensure that $\lim_{k\to\infty}E_{k}/\lambda_{k}=0.$ Combined with the fact that $\lim_{\lambda\to 0}\|x_{\lambda}-x_{\ast}\|_{\mathcal{X}}=0,$ this yields strong convergence $\lim_{k\to\infty}\|X_{k}-x_{\ast}\|_{\mathcal{X}}=0$ (both in $L^{2}$ and almost surely). Moreover, if a convergence rate for $\|x_{\lambda}-x_{\ast}\|_{\mathcal{X}}$ is known, the analysis allows us to also quantify a strong convergence rate for the iterates $X_{k}$ to $x_{\ast}$.